OpenStudy (mathisfun13):

Find the midpoint of the segment from point A (0.3,0.7) to point B (0.1,0.9) type the correct answer in (x,y) form

OpenStudy (anonymous):

formula is (x2+x1)/2 , (y2+y1)/2 in this case x1=0.3 x2=0.1 y1=0.7 y2=0.9 hence, (0.3+0.1)/2 , (0.9+0.7)/2

OpenStudy (anonymous):

i got (1/5,4/5)

OpenStudy (mathisfun13):

you use (x2-x1)2 +(y2-y1)2

OpenStudy (anonymous):

nono..(x2-x1)2 +(y2-y1)2 is used to find distance... (x2+x1)/2 , (y2+y1)/2 is for midpoint :)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Midpoint hope this help :)

ganeshie8 (ganeshie8):

A (0.3,0.7) to point B (0.1,0.9) x1 y1 x2 y2

OpenStudy (mathisfun13):

i got (-0.2)2 + (0.2)2

ganeshie8 (ganeshie8):

Midpoint = \((\frac{x1+x2}{2}, \frac{y1+y2}{2})\) = \((\frac{0.3 + 0.1}{2}, \frac{0.7 + 0.9}{2})\) = ?

OpenStudy (mathisfun13):

(0.4,1.6)?

ganeshie8 (ganeshie8):

dont forget to divide each wid 2

OpenStudy (mathisfun13):

(.2,.8)

ganeshie8 (ganeshie8):

\((\frac{0.4}{2}, \frac{1.6}{2})\)

ganeshie8 (ganeshie8):

Yes !!